OpenStudy (zela101):

Evaluate the integral: The parabolic arc y=x^2 from (0,0) to (1,1)

OpenStudy (zela101):

\[\int\limits_{c}^{}(3x+2y)dx+(2x-y)dy\]

OpenStudy (zela101):

How would i determine the t in x and y?

OpenStudy (irishboy123):

why do you have to \(y = x^2\) \(dy = 2x \, dx\)

OpenStudy (irishboy123):

\[\int\limits_{c}^{}(3x+2x^2)dx+(2x-x^2)2x dx\]

OpenStudy (zela101):

It makes sense from that perspective, but i want to replace everything by t.

OpenStudy (zela101):

@dan815

OpenStudy (zela101):

The book says x=t and y=x^2 

OpenStudy (dan815):